OpenStudy (anonymous):

i need help with 2x^2-x-6= 0 and y^2-6y+2=0 these are solving quadratic equations and i need the help quick please! need an easy explanation and i will give medal

OpenStudy (anonymous):

Chill... I'll help u..

ganeshie8 (ganeshie8):

which method you wanto use

ganeshie8 (ganeshie8):

1. factoring 2. completing square 3. quadratic formula

ganeshie8 (ganeshie8):

pick one

ganeshie8 (ganeshie8):

pick one that you're familiar wid already

ganeshie8 (ganeshie8):

ok you have 2 questions here. first question we will do by factoring. second we wil try completing square. okay ?

ganeshie8 (ganeshie8):

all 3 methods will give u same answer

ganeshie8 (ganeshie8):

2x^2-x-6= 0

ganeshie8 (ganeshie8):

-6*2 = -12 so we need to pick two numebers that multiply to -12, and addup to -1 (cuz middle term is -x)

ganeshie8 (ganeshie8):

can you pick the two numbers ?

ganeshie8 (ganeshie8):

will -4 and 3 work ?

ganeshie8 (ganeshie8):

lets see :- -4*3 = -12 -4+3 = -1

ganeshie8 (ganeshie8):

exactly ! so these are our golden numbers for this problem

ganeshie8 (ganeshie8):

2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0

ganeshie8 (ganeshie8):

i wrote -x as, -4x + 3x, which is legitimate

ganeshie8 (ganeshie8):

2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0

ganeshie8 (ganeshie8):

tell me whats common in first two terms

ganeshie8 (ganeshie8):

i can take out 2x, eh ?

ganeshie8 (ganeshie8):

2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0 2x(x-2) + 3x - 6 = 0

ganeshie8 (ganeshie8):

tell me whats common in last two terms ?

ganeshie8 (ganeshie8):

last two terms

ganeshie8 (ganeshie8):

2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0 2x(x-2) + 3x - 6 = 0 ^ ^

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

so take that out

ganeshie8 (ganeshie8):

correct ! 2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0 2x(x-2) + 3x - 6 = 0 2x(x-2) + 3(x-2) = 0

ganeshie8 (ganeshie8):

now do we see, (x-2) common in both terms ?

ganeshie8 (ganeshie8):

take that all the way out

ganeshie8 (ganeshie8):

2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0 2x(x-2) + 3x - 6 = 0 2x(x-2) + 3(x-2) = 0 (x-2)(2x+3) = 0

ganeshie8 (ganeshie8):

we factored fully ! now to the answer

ganeshie8 (ganeshie8):

2x^2-x-6= 0 2x^2 -4x + 3x -6 = 0 2x(x-2) + 3x - 6 = 0 2x(x-2) + 3(x-2) = 0 (x-2)(2x+3) = 0 x-2 = 0, or 2x+3 = 0 x = 2, or 2x = -3 x = 2, or x = -3/2

ganeshie8 (ganeshie8):

those are the solutions of x

ganeshie8 (ganeshie8):

makes some sense ?

ganeshie8 (ganeshie8):

we are done.

OpenStudy (anonymous):

i thought you were going to help me with both .. @ganeshie8

ganeshie8 (ganeshie8):

oh ya !!!

ganeshie8 (ganeshie8):

y^2-6y+2=0

OpenStudy (anonymous):

continue brb

ganeshie8 (ganeshie8):

we will solve this by completing the square ok

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

you can go, finish ur work, and come back. il be around :D

OpenStudy (anonymous):

oh i had to look at a reply i got lol i need help

ganeshie8 (ganeshie8):

hmm..

OpenStudy (anonymous):

people are still replying on my closed questions but im doing my work which is why i need help lol

OpenStudy (anonymous):

@ganeshie8 so completing the square?